# Generative AI for Intent-Driven Network Management in 6G: A Case Study on Hierarchical Learning Approach

### v1

@@ V Conclusions
This paper introduces a hierarchical GenAI-driven framework for IDN management in 6G environments by integrating GenAI across intent processing, validation, and execution. By fine-tuning an LLM with QLoRA for efficient intent processing, using transformer-based predictors for validation, and deploying the HDMGA for execution, the proposed approach enhances network automation. The case study demonstrated HDMGA’s superior performance in achieving network objectives like delay reduction, increased energy efficiency, and throughput while reducing computational overhead. In our future works, we wish to integrate RLHF for refining intent validation and execution processes to enable real-time learning and adaptation based on operator feedback and network conditions.
@@ Acknowledgment
This work has been supported by MITACS and Ericsson Canada, and NSERC Canada Research Chairs Program.

### v2

@@ V Conclusions
This article presented a GenAI-driven framework for IDN management that integrates selective state-space modeling across intent processing, validation, and execution for the first time in the literature. By utilizing a fine-tuned Mamba-based LLM for intent interpretation, a Mamba-enabled predictor for proactive validation, and a Mamba-enhanced intent execution module, the proposed architecture delivers a unified and scalable approach to intelligent automation in 6G RAN environments. The case study demonstrated that this Mamba’s SSM-driven design significantly improves delay, throughput, energy efficiency, and reliability compared with existing IDN management approaches, while reducing computational overhead through efficient sequence modeling. Looking ahead, scaling the proposed architecture to larger deployments or conducting prototype-level validation (e.g., on Open RAN compliant testbeds) represents an important direction for future work.
@@ Acknowledgment
This work has been supported by MITACS, Ericsson Canada, and NSERC Canada Research Chairs Program.